OpenStudy (anonymous):

A circular jogging track is being designed so that one complete lap around the track is one mile long. What should the radius of the track be to the nearest foot? (one mile equals five thousand two hundred eighty feet)

OpenStudy (anonymous):

5280ft = 1 mile

OpenStudy (anonymous):

ohh wait lol

OpenStudy (anonymous):

sorry one sec. :)

OpenStudy (anonymous):

ok

OpenStudy (kropot72):

\[circumference=2\times \pi \times radius\] \[radius=\frac{circumference}{2\times \pi}=\frac{5280}{2\times \pi}\]

OpenStudy (anonymous):

well I think that answered ur question.. :P @Isabella_Kapadia @kropot72